Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
Six years ago
Bob = 4Mary - 6
In four years
Bob = 2Mary + 4
4M - 6 = 2M + 4
2M = 10
Mary = 5 years old
Substitute both equations with 5, and Bob is 14
I just want to know if that is right or wrong?
: Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
Six years ago
Bob = 4Mary - 6
In four years
Bob = 2Mary + 4
4M - 6 = 2M + 4
2M = 10
Mary = 5 years old
Substitute both equations with 5, and Bob is 14
I just want to know if that is right or wrong?

Answer by oscargut(667)

Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
B= Bob's age now
M= Mary's age now
You did it wrong you have to start like this:
B-6=4(M-6)
B+4=2(M+4)
solving we have
B=26 and M=11
Question 151380: Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
Six years ago
Bob = 4Mary - 6
In four years
Bob = 2Mary + 4
4M - 6 = 2M + 4
2M = 10
Mary = 5 years old
Substitute both equations with 5, and Bob is 14
I just want to know if that is right or wrong?
: Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
Six years ago
Bob = 4Mary - 6
In four years
Bob = 2Mary + 4
4M - 6 = 2M + 4
2M = 10
Mary = 5 years old
Substitute both equations with 5, and Bob is 14
I just want to know if that is right or wrong?

Answer by stanbon(18792)

Six years ago bob was four times as old as Mary.
In four years he will be twice as old as Mary.
How old are they?
-----------
Let Mary's age now be "m".
Let Bob's age now be "b".
---------------------------
6 yrs. ago equation:
(b-6) = 4(m-6)
------------------------------
4 yrs. from now equation:
b = 2m
--------------------------
Substitute to solve for "m".
(2m-6) = 4(m-6)
2m-6 = 4m-24
2m = 18
m = 9 yrs old (Mary's age now)
-------------------
Since b = 2m, b = 2*9 = 18 yrs old (Bob's age now)
===========================
Cheers,
Stan H.